OpenStudy (anonymous):

1. Describe the relationship between compound interest and exponential growth. 2. Describe how the account balance of a savings account earning simple interest is determined. 3. Explain the difference between a savings account earning simple interest and a savings account earning compound interest. 4. Which type of savings account would you open and why? Please help.

OpenStudy (cwrw238):

Here is an example of compound interest 100 dollars invested for 3 years at 10% per year amount after 3 years = 100(1 + 0.10)^3 = 133.10 Now suppose interest is paid every 6 months instead of yearly that means that there will be 6 payments over the 3 years of 5% amount after 3 years = 100(1 + 0.05)^6 = 134.01 so the more often the interest is paid the more you get no if you g to the extreme and pay evry minute you can expect a lot more! every second - a lot lot more the limit of this will be when n (the number of times paid) approaches infinity the exponential constant e = limit n--> infinity of (1 + 1/n)^n its values is 2.71828 to 5 places of decimals try working out (1 + 1/1000)^1000 on your calculator - yo'll get a value quite close to the constant e

OpenStudy (anonymous):

Im still kinda confused....

OpenStudy (cwrw238):

maybe if I used fractions instead of decimals for compound interest for 3 years annual payments 1 % = 1/100 = 0.01 we have ( 1 + 0.01)^3 -3 payments payments every 6 months in terst will be half of 10 % = 0.10/2 we have (1 + 0.01/2)^(3*2) for 4 payments a year we have ( 1 + 0.01/4)^(3*4) for n years its (1 + 1/n)^n now if n becomes a very big number the value of (1 + 1/n)^n will approach a limit which is about 2.71828 which is assigned the number e - the exponential constant

OpenStudy (anonymous):

I just need a description, not really any examples.

OpenStudy (cwrw238):

so if any bank offer you exponential compound interest - grab it - you'll soon be very rich.

OpenStudy (cwrw238):

its hard to explain without giving examples

OpenStudy (anonymous):

Haha thanks for the advice :) I know, its just I only need a description :/

OpenStudy (tkhunny):

They are the same. Compound interest is modeled and calculated by exponential growth.

OpenStudy (republican):

tkhunny they are not the same.

OpenStudy (tkhunny):

I repeat, They are the same. Compound interest is modeled and calculated by exponential growth.